# Revisiting Quantization of Gauge Field Theories: Sandwich Quantization Scheme

### v2

@@
Quantization of field theories with gauge symmetry is an extensively discussed and well-established topic. In this short note, we revisit this old problem. The gauge degrees of freedom have vanishing momenta, and hence their equations of motion appear as constraints on the system. We argue that to ensure consistency of quantization one can impose these constraints as “sandwich conditions”: The physical Hilbert space of the theory consists of all states for which the constraints sandwiched between any two physical states vanish. We solve the sandwich constraints and show they have solutions not discussed in the gauge field theory literature. We briefly discuss the physical meaning of these solutions and implications of the sandwich quantization scheme.
@@ 1 Introduction
In general, one may view particle theory as a $0+1$ dimensional field theory residing on the particle worldline. Our discussions here hold for a generic $d+1$ dimensional field theory for any $d$, including $d=0$. Thus, in what follows when we say (quantum) field theory, it includes the particle theory as well. In canonical quantization procedure, we should also define Hilbert space of the theory over which the operators act. The Hilbert space is usually constructed based on a “vacuum state” and all the other states in the Hilbert space of a (local) QFT are constructed by the action of local operators on the vacuum state [1, 2]. Therefore, by construction, we have operator-state correspondence. [^1] We note that the operator-state correspondence introduced here and we use in this paper need not be a one-to-one correspondence (as is more commonly used and stated in 2d Conformal Field Theories). For our purpose it is enough if the correspondence is onto, i.e. all states in the Hilbert space be constructed by the action of at least one operator on a vacuum state. We thank Glenn Barnich and Marco Serone for a comment on this point. (Vacuum state then corresponds to the identity operator.)
@@
Besides the canonical quantization, there is the well-established and commonly used path integral formulation. [^2] This is despite the fact that a rigorous mathematical definition of the path integral and in particular its measure, is still missing. There is an extensive literature establishing the canonical quantization and path integral quantization schemes agree on all physical observables. In the path integral formulation of local QFTs we do not have Hilbert space and the path integral computes all observables, that are assumed to be of the form of generic $n$-point functions, vacuum expectation value (VEV) of generic products of local operators [1, 2]. Nevertheless, one can implicitly talk/think about states and Hilbert space of the theory assuming operator-state correspondence (see footnote 1).
@@
Gauge theories and gauge symmetries have been the cornerstone of physical formulations in the last century. Gauge symmetries, despite the commonly used name, do not point to conserved charges via Noether’s first theorem [1, 3]. They are rather “redundancies of description” usually introduced to make other symmetries of the system more manifest or as a theoretical guiding principle for fixing interactions among fields. In particle theory, the gauge symmetry corresponds to reparametrization of the worldline parameter. In the context of field theories, gauge symmetries may correspond to “internal symmetries” over the field space, like the case of Maxwell theory or non-Abelian gauge theories in the standard model of particle physics, or to “external (spacetime)” symmetries, like diffeomorphism invariance in gravitational theories [4].
@@
Gauge theories are prime examples of constrained systems: Some of the degrees of freedom (d.o.f) in the Lagrangian of gauge theories, gauge d.o.f, have vanishing conjugate momenta. Thus, equations of motion for the gauge d.o.f. are constraints and gauge d.o.f are not propagating and dynamical (in contrast to a dynamical equation involving second order time derivatives). As such, systems with gauge symmetry are constrained systems and Dirac’s procedure [5] or other suitable methods for quantizing constrained systems should be invoked [6].
@@
The standard procedure for dealing with gauge field theories starts with gauge-fixing: If we have $N$ gauge d.o.f, gauge-fixing involves assuming $N$ arbitrary relations among gauge fields and possibly their derivative such that these relations under a generic gauge transformation do not remain invariant and hence fixing these relations amounts to fixing a gauge. Gauge-fixing relations may be viewed as constraints which together with the equations of motion (EoM) for the gauge d.o.f form a well-posed constraint system [5, 6]. [^3] Note that while the gauge-fixing relations, by construction and definition, are not gauge invariant, equations of motion for the gauge d.o.f, by the very definition of gauge theory, are gauge covariant: the set of EoM is closed under gauge transformations. One may then construct solution space of the classical gauge-fixed theory and go through the standard canonical quantization procedure. Alternatively, one may choose the path integral method.
@@
In this work, we revisit quantization of gauge field theories, especially the canonical quantization. After a quick but careful review of the standard discussions, we observe that EoM of the gauge d.o.f. need not necessarily be imposed as one finds in the standard textbooks; one can impose them as “sandwich conditions”. That is, physical Hilbert space of the gauge field theory may be defined upon the condition that the EoM of gauge d.o.f, promoted to operators upon quantization, vanish when sandwiched between any two physical states.
2 Gauge theory, classic treatment
@@
    Gauge d.o.f $\varphi_{i}$, by definition, are the fields with identically vanishing conjugate momentum. If $t$ denotes the time direction, that is,
@@
${\cal G}_{i}(\varphi_{i},\psi_{A})=0,$ (4)
where ${\cal G}_{i}$, which specify the desired gauge-fixing of our choice, is a function(al) not invariant under gauge transformations. In other words, to have a consistent gauge-fixing $\det\left(\frac{\delta{\cal G}_{i}}{\delta\lambda_{j}}\right)\neq 0$. As such, (4) fixes the freedom in choices of $\lambda_{i}$: under a gauge transformation ${\cal G}_{i}=0$ will not hold and requiring (4) amounts to specific choice for $\lambda_{i}$.
Consistency of the gauge-fixing requires that (4) is compatible with the evolution of the system [6]; it should hold at all times, in particular ${\cal G}_{i}=0$ and ${\cal C}^{i}=0$ should hold simultaneously. This consistency requirement may be analyzed and carried out in the Hamiltonian formulation, yielding infinite chain of secondary constraints [6] or in the Lagrangian (action) formulation [1, 2, 9] with physically identical results. A particularly simple and handy gauge-fixing, that we adopt in the rest of this work, is to fix the functional form of gauge d.of. as
@@
where $\varphi^{0}_{i}$ is a given function with no dependence on other fields; temporal or axial gauge-fixing is an example of this choice. The classically physical (gauge inequivalent) field configurations are hence configurations of $\psi_{A}$ subject to their EoM with $\varphi_{i}-\varphi_{i}^{0}=0$ and ${\cal C}^{i}=0$ (at all times). As we will discuss in more detail in section 4, for the case of Maxwell theory this physical solution space is described by transverse polarization of electromagnetic waves (with $d-2$ polarizations for the $d$ dimensional theory) plus the boundary soft modes specified by functions on codimension 2 celestial sphere. See [10] for more details.
Right-Action Quantization Scheme.
@@
Having the classical physical solution space, one may proceed with the canonical quantization, i.e. one may promote the physical field configurations and their momenta to operators that satisfy the canonical commutation relations. This guarantees the expectation that one-particle Hilbert space of the quantum theory and the classical field configurations are in one-to-one relation. This is the standard procedure introduced by Julian Schwinger in 1948 [11, 12] and developed further in 1950 by Gupta [13] and Bleuler [14]. See [9] for a review of the Gupta-Bleuler method which is a well-known but old and somewhat outdated quantization method. The Gupta-Bleuler method was surpassed by the BRST method [15].
In what follows, the operator associated with a generic function(al) of fields ${\cal O}[\Phi]$ will be denoted by $\bm{\mathcal{O}}$, the Hilbert space of all field configurations by $\mathcal{H}_{\text{t}}$ and the Hilbert space of all physical field configurations by $\mathcal{H}_{\text{p}}$. One can construct $\mathcal{H}_{\text{t}}$ as in standard QFT textbooks: construct solutions to the equations of motion of classical EoM and promote the “Fourier coefficients” to canonical creation/annihilation operators or positive and negative frequency modes. One-particle sector of $\mathcal{H}_{\text{t}}$ is then constructed by the action of the creation operators on the vacuum state $|0\rangle$ and the whole $\mathcal{H}_{\text{t}}$ is the Fock space built upon this one-particle Hilbert space. In this notation, functionals associated with (5) and the EoM of $\varphi_{i}$ will be denoted by $\bm{\varphi}_{i}-\bm{\varphi}^{0}_{i}$ and $\bm{{\cal C}}^{i}$, respectively and the positive frequency sector of these will be denoted by $\left(\bm{\varphi}_{i}-\bm{\varphi}^{0}_{i}\right)^{+}$ and $\left(\bm{{\cal C}}^{i}\right)^{+}$.
@@
When we discuss vanishing of an operator we should specify over which Hilbert space it vanishes. In our case, $\bm{\varphi}_{i}-\bm{\varphi}^{0}_{i}$ and $\bm{{\cal C}}^{i}$ (or the positive frequency parts of them) do not vanish over $\mathcal{H}_{\text{t}}$ but are zero over $\mathcal{H}_{\text{p}}$. To be precise, the Schwinger-Gupta-Bleuler quantization scheme [11, 12, 13, 14], or the “right-action quantization scheme” requires [^4] Since the positive and negative frequency parts of an operator by construction do not commute with each other, it is clear that one cannot require $\left(\bm{\varphi}_{i}-\bm{\varphi}^{0}_{i}\right)^{+}|\psi\rangle=0$ and $\left(\bm{\varphi}_{i}-\bm{\varphi}^{0}_{i}\right)^{-}|\psi\rangle=0$ simultaneously.
@@ 3 Sandwich quantization scheme
We start with recalling that operators on a Hilbert space $\mathcal{H}_{\text{t}}$ may customarily be viewed as matrices over the Hilbert space, i.e. a given operator $\bm{{\cal O}}$ over $\mathcal{H}_{\text{t}}$ may be replaced with $\langle\psi|\bm{\mathcal{O}}|\phi\rangle$ for any two states $|\phi\rangle,|\psi\rangle\in\mathcal{H}_{\text{t}}$. So, a natural choice/proposal to replace vanishing of an operator by vanishing of its matrix elements. Explicitly, one may explore replacing (6) with the “sandwich constraints”:
@@
One can adopt a quantization scheme with a mixture of sandwich and right-action conditions are used: impose gauge-fixing condition as a right-action (impose it at classical level) and impose $\bm{\mathcal{C}}^{i}$ as a sandwich constraint. We adopt this quantization scheme, explicitly:
${\varphi}_{i}-{\varphi}^{0}_{i}=0,\ \ \langle\psi^{\prime}|\bm{\mathcal{C}}^{i}|\psi\rangle=0,\ \ \forall|\psi\rangle,|\psi^{\prime}\rangle\in\mathcal{H}_{\text{p}}$ (8)
That is, we construct the classical solution space obtained in the (axial-type) ${\varphi}_{i}-{\varphi}^{0}_{i}=0$ gauge. Then, quantize this and construct the $\mathcal{H}_{\text{t}}$ by promoting these solutions to operators acting on vacuum state (eliminated by annihilation operators). The physical Hilbert space $\mathcal{H}_{\text{p}}$ is then a subset of $\mathcal{H}_{\text{t}}$ that satisfies (8). We argue in the next subsection that the above sandwich quantization scheme (8) may be obtained via path integral quantization.
We close this part by the remark that it was already noted in the old literature [13, 14] (see also [16, 17]) that (7) are physically what one needs for quantization. However, in the Gupta-Bleuler quantization only the gauge-fixing condition was required to hold as the sandwich condition (the left equation in (7)) and in practice it was imposed and solved by the right-action requirement; the EoM for the gauge d.o.f $\bm{\mathcal{C}}^{i}=0$ was not considered.
@@ 3.1 Path integral derivation of sandwich quantization scheme (8)
Physical observables are then generic $n$-point functions of these operators, VEV of time-ordered product of these operators: [^5] We are assuming that vacuum state $|0\rangle$ is a state in $\mathcal{H}_{\text{t}}$.
@@
$\begin{split}\langle\bm{\mathcal{O}}_{1}(x_{1})\bm{\mathcal{O}}_{2}(x_{2})\ {\bm{\mathcal{C}}}^{i}(x)\cdots\bm{\mathcal{O}}_{n}(x_{n})\rangle&=\int\overline{{\cal D}\Phi}\ {\cal O}_{1}(x_{1}){\cal O}_{2}(x_{2})\cdots{\cal O}_{n}(x_{n})\ {\cal C}^{i}(x)\ e^{\frac{i}{\hbar}\bar{S}[\Phi]}\\ &=-i{\hbar}\int\overline{{\cal D}\Phi}\ {\cal O}_{1}(x_{1}){\cal O}_{2}(x_{2})\cdots{\cal O}_{n}(x_{n})\ \left(\frac{\delta}{\delta\varphi_{i}(x)}e^{\frac{i}{\hbar}S[\Phi]}\right)_{\varphi_{i}=\varphi_{i}^{0}}\\ &=i{\hbar}\int\overline{{\cal D}\Phi}\ \frac{\delta}{\delta\varphi_{i}(x)}\left({\cal O}_{1}(x_{1}){\cal O}_{2}(x_{2})\cdots{\cal O}_{n}(x_{n})\right)\big{|}_{\varphi_{i}=\varphi_{i}^{0}}\ e^{\frac{i}{\hbar}\bar{S}[\Phi]}\end{split}$ (12)
@@
Noting that ${\bm{\mathcal{O}}}_{i}$ are gauge-invariant operators (9), we learn that
$\langle\bm{\mathcal{O}}_{1}(x_{1})\cdots\ {\bm{\mathcal{C}}}^{i}(x)\cdots\bm{\mathcal{O}}_{n}(x_{n})\rangle=0\qquad\forall\ \text{local gauge-invariant operators}\ {\bm{\mathcal{O}}}_{i}.$ (14)
One may recall operator-state correspondence (see footnote 1), that to any local gauge-invariant operator ${\bm{\mathcal{O}}}_{i}$ one may associate a state, conveniently denoted by $|{\cal O}_{i}\rangle$, the vacuum state $|0\rangle$ corresponds to the identity operator and a generic (multi-particle) state may be associated with products of such operators. In this terminology, (14) is equivalent to sandwich constraints (8). In other words, we have presented a derivation of the sandwich quantization scheme (8) from the path integral formulation. This derivation also implies that if $\bm{\mathcal{O}}_{1}$ and $\bm{\mathcal{O}}_{2}$ are two physical operators ($|{\cal O}_{1}\rangle,|{\cal O}_{2}\rangle\in\mathcal{H}_{\text{p}}$), then $\bm{\mathcal{O}}_{1}\bm{\mathcal{O}}_{2}|0\rangle$ should also be in $\mathcal{H}_{\text{p}}$; see appendix B of [19] for related arguments.
Before moving on and discussing the proposal for solving the sandwich constraints, we remark that the right-hand-side of (13) is proportional to $\hbar$. At the quantum level one may have defined the constraints as $\frac{1}{\hbar}\bm{\mathcal{C}}^{i}$, making more manifest that the sandwich quantization scheme is an option appearing at the quantum level, while at the classical level we deal with ${\cal C}^{i}=0$.

### v3

@@
Quantization of field theories with gauge symmetry is an extensively discussed and well-established topic. In this short note, we revisit this old problem. While we confirm all details of the existing literature, we highlight a potentially important point which may provide a better understanding of and insights on the quantization of gauged systems. The gauge degrees of freedom have vanishing momenta, and hence their equations of motion appear as constraints on the system. We argue that to ensure consistency of quantization one can impose these constraints as “sandwich conditions”: The physical Hilbert space of the theory consists of all states for which the constraints sandwiched between any two physical states vanish. We solve the sandwich constraints and show they have solutions not discussed in the gauge field theory literature. We briefly discuss the physical meaning of these solutions and implications of the sandwich quantization scheme.
@@ 1 Introduction
In general, one may view particle theory as a $0+1$ dimensional field theory residing on the particle worldline. Our discussions here hold for a generic $d+1$ dimensional field theory for any $d$, including $d=0$. Thus, in what follows when we say (quantum) field theory, it includes the particle theory as well. In canonical quantization procedure, we should also define Hilbert space of the theory over which the operators act. The Hilbert space is usually constructed based on a “vacuum state” and all the other states in the Hilbert space of a (local) QFT are constructed by the action of local operators on the vacuum state [23, 14]. Therefore, by construction, we have a version of operator-state correspondence (OSC). We note that the OSC introduced here and we use in this paper need not be a one-to-one correspondence (as is more commonly used and stated in 2d Conformal Field Theories, e.g. see [15]). What we mean by OSC is that all states in the Hilbert space are constructed by the action of at least one operator on a vacuum state. [^1] We thank Glenn Barnich and Marco Serone for a comment on this point. (Vacuum state then corresponds to the identity operator.)
@@
Besides the canonical quantization, there is the well-established and commonly used path integral formulation. [^2] This is despite the fact that a rigorous mathematical definition of the path integral and in particular its measure, is still missing. There is an extensive literature establishing the canonical quantization and path integral quantization schemes agree on all physical observables. In the path integral formulation of local QFTs we do not have the Hilbert space and the path integral computes all physical observables, that are assumed to be of the form of generic $n$-point functions, vacuum expectation value (VEV) of generic products of local operators [23, 14]. Nevertheless, one can implicitly talk/think about states and Hilbert space of the theory assuming the OSC defined above.
@@
Gauge theories and gauge symmetries have been the cornerstone of physical formulations in the last century. Gauge symmetries, despite the commonly used name, do not point to conserved charges via Noether’s first theorem [23, 24]. They are rather “redundancies of description” usually introduced to make other symmetries of the system more manifest or as a theoretical guiding principle for fixing interactions among fields. In particle theory, the gauge symmetry corresponds to reparameterization of the worldline parameter. In the context of field theories, gauge symmetries may correspond to “internal symmetries” over the field space, like the case of Maxwell theory or non-Abelian gauge theories in the standard model of particle physics, or to “external (spacetime)” symmetries, like diffeomorphism invariance in gravitational theories [22].
@@
Gauge theories are prime examples of constrained systems: Some of the degrees of freedom (d.o.f) in the Lagrangian of gauge theories, the gauge d.o.f, have vanishing conjugate momenta. Thus, equations of motion for the gauge d.o.f. are constraints and gauge d.o.f are not propagating and dynamical (in contrast to a dynamical equation involving second order time derivatives). As such, systems with gauge symmetry are constrained systems and Dirac’s procedure [7] or other suitable methods for quantizing constrained systems should be invoked [10].
@@
The standard procedure for dealing with gauge field theories starts with gauge-fixing: If we have $N$ gauge d.o.f, gauge-fixing involves assuming $N$ arbitrary relations among gauge fields and possibly their derivative such that these relations do not remain invariant under a generic gauge transformation. Thus, fixing these relations amounts to fixing a gauge. Gauge-fixing relations may be viewed as constraints which together with the equations of motion (EoM) for the gauge d.o.f form a well-posed constraint system [7, 10]. [^3] Note that while the gauge-fixing relations, by construction and definition, are not gauge invariant, equations of motion for the gauge d.o.f, by the very definition of gauge theory, are gauge covariant: the set of EoM is closed under gauge transformations. One may then construct solution space of the classical gauge-fixed theory and go through the standard canonical quantization procedure. Alternatively, one may choose the path integral method.
@@
In this note, we revisit quantization of gauge field theories, especially the canonical quantization. We start with a quick but careful review of the standard discussions and observe that EoM of the gauge d.o.f. need not necessarily be imposed as one finds in the standard textbooks; one can impose them as “sandwich conditions”. That is, physical Hilbert space of the gauge field theory may be defined upon the condition that the EoM of gauge d.o.f, promoted to operators upon quantization, vanish when sandwiched between any two physical states.
Disclaimer: We stress that this note is prepared for a special issue trying to stress simple and basic facts and issues of educational value regarding quantization of gauged systems. By no means, we do not intend to provide a full account of the problem, our goal is to highlight a point, a concept, that has not been noted before and can shed a better light on the gauge systems with potentially interesting and important physical significance, yet to be developed and uncovered.
2 Gauge theories, a quick review of basic facts
@@
    Gauge d.o.f $\varphi_{i}$, by definition, are the fields with identically vanishing conjugate momentum. If $t$ denotes the time direction, that is, [^4] One should note that while all the gauge d.o.f have vanishing conjugate momentum, the converse is not necessarily true. For example, consider the Maxwell theory plus a mass term for the vector field, where $A_{0}$ has a vanishing momentum (cf. section 4), this theory has no gauge symmetry. We thank the anonymous referee for pointing this out.
@@ 2.1 Gauge-fixing
${\cal G}_{i}(\varphi_{i},\psi_{A})=0\,.$ (4)
${\cal G}_{i}$ specify the desired gauge-fixing of our choice and are functionals not invariant under gauge transformations. In other words, to have a consistent gauge-fixing $\det\left(\frac{\delta{\cal G}_{i}}{\delta\lambda_{j}}\right)\neq 0$ where $\frac{\delta{\cal G}_{i}}{\delta\lambda_{j}}=\frac{\delta{\cal G}_{i}}{\delta\Phi}{\delta}_{\lambda_{j}}\Phi$}. As such, (4) fixes the freedom in choices of $\lambda_{i}$: under a gauge transformation ${\cal G}_{i}=0$ will not hold and requiring (4) amounts to specific choice for $\lambda_{i}$.
Consistency of the gauge-fixing requires that (4) is compatible with the evolution of the system [10], i.e. it should hold at all times, in particular, ${\cal G}_{i}=0$ and ${\cal C}^{i}=0$ should hold simultaneously. This consistency requirement may be analyzed and carried out in the Hamiltonian formulation, yielding the chain of secondary constraints [10] or, in the Lagrangian (action) formulation [23, 14, 11], yielding physically identical results. A particularly simple and handy gauge-fixing, that we adopt in the rest of this work, is to fix the functional form of gauge d.of. as
@@
where $\varphi^{0}_{i}$ is a given function with no dependence on other fields; temporal or axial gauge-fixing is an example of the gauge-fixing in (5).
@@
The classically physical (gauge inequivalent) field configurations are hence configurations of $\psi_{A}$ subject to their EoM with $\varphi_{i}-\varphi_{i}^{0}=0$ and ${\cal C}^{i}=0$ (at all times). As we will discuss in more detail in section 4, for the case of Maxwell theory this physical solution space is described by transverse polarization of electromagnetic waves (with d2d-2 polarizations for the $d$ dimensional theory) plus the boundary soft modes specified by functions on codimension 2 celestial sphere. See [20] for more details.
2.2 Right-action quantization scheme
@@
Having the classical physical solution space, one may proceed with the canonical quantization, i.e. one may promote the physical field configurations and their momenta to operators that satisfy the canonical commutation relations. This guarantees the expectation that one-particle Hilbert space of the quantum theory and the classical field configurations are in one-to-one relation. This is the standard procedure introduced by Julian Schwinger in 1948 [17, 18] and developed further in 1950 by Gupta [9] and Bleuler [5]. See [11] for a review of the Gupta-Bleuler method which is a well-known but old and somewhat outdated quantization method. The Gupta-Bleuler method was surpassed by the BRST method [10, 4, 21].
In what follows, the operator associated with a generic function(al) of fields ${\cal O}[\Phi]$ will be denoted by $\bm{\mathcal{O}}$, the vector space of all field configurations by $\mathcal{H}_{\text{tot}}$ and the Hilbert space of all physical field configurations by $\mathcal{H}_{\text{p}}$. One can construct $\mathcal{H}_{\text{tot}}$ as in standard QFT textbooks:
  1. 1.
    Construct solutions to the classical EoM;
  2. 2.
    Promote the “Fourier coefficients” of the solutions to the canonical creation/annihilation operators or to the positive and negative frequency modes.
  3. 3.
    Define the vacuum state $|0\rangle$ as the state annihilated by all annihilation operators.
  4. 4.
    Construct one-particle sector of $\mathcal{H}_{\text{tot}}$ by the action of the creation operators on the vacuum state $|0\rangle$. The whole $\mathcal{H}_{\text{tot}}$ is the Fock space built upon this one-particle vector space.
  5. 5.
    Note that for gauge theories, generically, $\mathcal{H}_{\text{tot}}$ is not a Hilbert space, because it contains zero or negative norm states. After removing these states (using ghosts or other methods), we obtain the physical Hilbert space $\mathcal{H}_{\text{p}}$ which only contains positive and finite norm states; physical states may be normalized to have norm 1.
The main point of the gauge-fixing procedure is item 5. above and removing the unphysical states/d.o.f. In the now-standard textbook treatment, removal of these states is performed through ghosts and is guaranteed by the BRST symmetry [14, 24, 10, 11]. Here, we will not take the route of ghosts and the BRST; for illustrative purposes of our main idea, we follow axial gauge-fixing in which it is known that the ghosts decouple and we revisit Gupta-Bleuler procedure.
@@
In the adopted notation, functionals associated with (5) and the EoM of $\varphi_{i}$ will be respectively denoted by $\bm{\varphi}_{i}-\bm{\varphi}^{0}_{i}$ and $\bm{{\cal C}}^{i}$ and, the positive frequency sector of these, the part consisting of creation operators, will be denoted by $\left(\bm{\varphi}_{i}-\bm{\varphi}^{0}_{i}\right)^{+}$ and $\left(\bm{{\cal C}}^{i}\right)^{+}$. When we discuss vanishing of an operator we should specify over which vector/Hilbert space it vanishes. In our case, $\bm{\varphi}_{i}-\bm{\varphi}^{0}_{i}$ and $\bm{{\cal C}}^{i}$ (or the positive frequency parts of them) do not vanish over $\mathcal{H}_{\text{tot}}$ but are zero over $\mathcal{H}_{\text{p}}$. To be precise, the Schwinger-Gupta-Bleuler quantization scheme [17, 18, 9, 5], or the “right-action quantization scheme” requires [^5] Since the positive and negative frequency parts of an operator by construction do not commute with each other, it is clear that one cannot require $\left(\bm{\varphi}_{i}-\bm{\varphi}^{0}_{i}\right)^{+}|\psi\rangle=0$ and $\left(\bm{\varphi}_{i}-\bm{\varphi}^{0}_{i}\right)^{-}|\psi\rangle=0$ simultaneously.
@@ 3 Sandwich quantization scheme
We start with recalling that operators on a vector space $\mathcal{H}_{\text{tot}}$ may customarily be viewed as matrices over the vector space, i.e. a given operator $\bm{{\cal O}}$ over $\mathcal{H}_{\text{tot}}$ may be replaced with $\langle\psi|\bm{\mathcal{O}}|\phi\rangle$ for any two states $|\phi\rangle,|\psi\rangle\in\mathcal{H}_{\text{tot}}$. So, a natural choice/proposal is to replace vanishing of an operator by vanishing of its matrix elements. Explicitly, one may explore replacing (6) with the “sandwich constraints”:
@@
One can adopt a quantization scheme using a mixture of sandwich and right-action conditions: impose gauge-fixing condition at classical level and impose $\bm{\mathcal{C}}^{i}$ as a sandwich constraint. We adopt this quantization scheme, explicitly:
$\hskip-11.38109pt{\varphi}_{i}-{\varphi}^{0}_{i}=0,\ \ \langle\psi^{\prime}|\bm{\mathcal{C}}^{i}|\psi\rangle=0,\ \ \forall|\psi\rangle,|\psi^{\prime}\rangle\in\mathcal{H}_{\text{p}}\hskip-11.38109pt$ (8)
To proceed, one should first construct the classical solution space obtained in the (axial-type) ${\varphi}_{i}-{\varphi}^{0}_{i}=0$ gauge. Then, one should quantize the solution space and construct the $\mathcal{H}_{\text{tot}}$ by promoting these solutions to operators acting on vacuum state (which is defined as the state eliminated by annihilation operators). The physical Hilbert space $\mathcal{H}_{\text{p}}$ is then a subset of $\mathcal{H}_{\text{tot}}$ that satisfies (8). We argue in the next subsection that the above sandwich quantization scheme (8) may be obtained via path integral quantization.
We close this part by the remark that it was already noted in the old literature [9, 5] (see also [12, 13]) that (7) are what one physically needs for quantization. However, in the Gupta-Bleuler quantization only the gauge-fixing condition was required to hold as the sandwich condition (the left equation in (7)) and in practice, it was imposed and solved by the right-action requirement; the EoM for the gauge d.o.f $\bm{\mathcal{C}}^{i}=0$ was not considered.
@@ 3.1 Path integral derivation of sandwich quantization scheme (8)
Physical observables are then generic $n$-point functions of these operators, VEV of time-ordered product of these operators: [^6] We are assuming that vacuum state $|0\rangle$ is a state in $\mathcal{H}_{\text{tot}}$.
@@
$\begin{split}\hskip-14.22636pt\langle\bm{\mathcal{O}}_{1}(x_{1})\bm{\mathcal{O}}_{2}(x_{2})\ {\bm{\mathcal{C}}}^{i}(x)\cdots\bm{\mathcal{O}}_{n}(x_{n})\rangle&:=\int\overline{{\cal D}\Phi}\ {\cal O}_{1}(x_{1}){\cal O}_{2}(x_{2})\cdots{\cal O}_{n}(x_{n})\ {\cal C}^{i}(x)\ e^{\frac{i}{\hbar}\bar{S}[\Phi]}\\ &=-i{\hbar}\int\overline{{\cal D}\Phi}\ {\cal O}_{1}(x_{1}){\cal O}_{2}(x_{2})\cdots{\cal O}_{n}(x_{n})\ \left(\frac{\delta}{\delta\varphi_{i}(x)}e^{\frac{i}{\hbar}S[\Phi]}\right)_{\varphi_{i}=\varphi_{i}^{0}}\\ &=i{\hbar}\int\overline{{\cal D}\Phi}\ \frac{\delta}{\delta\varphi_{i}(x)}\left({\cal O}_{1}(x_{1}){\cal O}_{2}(x_{2})\cdots{\cal O}_{n}(x_{n})\right)\big|_{\varphi_{i}=\varphi_{i}^{0}}\ e^{\frac{i}{\hbar}\bar{S}[\Phi]}\end{split}$ (12)
@@
Noting that ${\bm{\mathcal{O}}}_{i}$ are gauge-invariant local operators (9), we learn that
$\hskip-11.38109pt\langle\bm{\mathcal{O}}_{1}(x_{1})\cdots\ {\bm{\mathcal{C}}}^{i}(x)\cdots\bm{\mathcal{O}}_{n}(x_{n})\rangle=0\qquad\forall\ \text{local gauge-invariant operators}\ {\bm{\mathcal{O}}}_{i}.\hskip-11.38109pt$ (14)
One may recall the OSC defined in the introduction, that to any local gauge-invariant operator ${\bm{\mathcal{O}}}_{i}$ one may associate a state, conveniently denoted by $|{\cal O}_{i}\rangle$. The vacuum state $|0\rangle$ corresponds to the identity operator and a generic (multi-particle) state may be associated with products of such operators. In this terminology, (14) is equivalent to sandwich constraints (8). In other words, we have presented a derivation of the sandwich quantization scheme (8) from the path integral formulation. This derivation also implies that if $\bm{\mathcal{O}}_{1}$ and $\bm{\mathcal{O}}_{2}$ are two physical operators ($|{\cal O}_{1}\rangle,|{\cal O}_{2}\rangle\in\mathcal{H}_{\text{p}}$), then $\bm{\mathcal{O}}_{1}\bm{\mathcal{O}}_{2}|0\rangle$ should also be in $\mathcal{H}_{\text{p}}$; see appendix B of [8] for related arguments.
Before moving on and discussing the proposal for solving the sandwich constraints, we remark that the right-hand-side of (13) is proportional to $\hbar$. At the quantum level, one may have defined the constraints as $\frac{1}{\hbar}\bm{\mathcal{C}}^{i}$, making more manifest that the sandwich quantization scheme is an option appearing at the quantum level, while at the classical level we deal with ${\cal C}^{i}=0$.